A calorimeter that contains 3.6 L of water is used to measure the enthalpy of solution for 45 g of KOH. Calculate the molar enthalpy of solution that is associated with the reaction if the temperature of the water changes from 25.5 °C to 20.7 °C. [The molar mass of KOH is 56.10 g/mol]  *


1
Expert's answer
2021-12-14T14:18:05-0500

Q = C x m x deltaT

deltaT = 25.5 - 20.7 = 4.8°C

Cwater = 4.186 J/g°C

Density of water is 1 g/L. Therefore, 3.6 L of water = 3.6 kg = 3600 g.

Q = 4.186 x 3600 x 4.8 = 72334 J


n (KOH) = m /M = 45 / 56.10 = 0.81 mol


Molar enthalpy of KOH solution = 72334 / 0.81 = 89 301 J/mol